OpenStudy (anonymous):

Complete this table of values for y=2x^2-4x-1 x = -2, -1, 0, 1, 2, 3, y = 5 -1 -1

OpenStudy (anonymous):

You just plug in the x values in the equation. y=2(-2)^2-4(2)-1 = 8-8-1=-1 y=2(1)^2-4(1)-1 = 2-4-1 = -3 y=2(3)^2-4(3)-1 = 18-12-1 = 5 So, (-2,-1) (1,-3) (3,5)